OpenStudy (anonymous):

************MEDALS************** 2(x-4)=4

OpenStudy (igreen):

First simplify the left side. \(2(x - 4) \rightarrow (2 \times x) + (2 \times -4)\) Can you simplify \((2 \times x) + (2 \times -4)\)? @The_Jokers_wife

OpenStudy (conqueror):

2(-4 + x) = 4 (-4 * 2 + x * 2) = 4 (-8 + 2x) = 4

OpenStudy (anonymous):

4 @iGreen

OpenStudy (conqueror):

Now solve it. Can you do that?

OpenStudy (igreen):

How did you get 4? \(2 \times x = 2x\) \(2 \times -4 = -8\) That gives you: \(2x - 8 = 4\) Add 8 to both sides: \(2x = 12\) Now can you divide 12 / 2? @The_Jokers_wife

OpenStudy (anonymous):

6 @iGreen

OpenStudy (igreen):

Yep, so that's your answer.